### v1

@@ 5.1. The Anosov property
Since $z\notin I$ there exist $n_{0}\in\mathbb{N}$ such that $q_{n_{0}}$ is positive on $z$. For $n\geq n_{0}$, let $z_{1},z_{2}$ be the two intersections of $L$ with the zeroes of $q_{n}$ so that the points $(x,z_{n}^{-},z,z_{n}^{+},y)$ are cyclically ordered. The sequences $(z_{n}^{-})$ and $(z_{n}^{+})$ are monotonic in $S$ and must converge to $x$ and $y$ since $S\cap I=\emptyset$. The value of $\text{cr}([q_{n_{0}}],[q_{n}])$ is bounded from above by the cross ratio $[z_{n_{0}}^{+},z_{n}^{+},z_{n}^{-},z_{n_{0}}^{-}]$, which is turns converges to the cross ratio $[z_{n_{0}}^{+},x,y,z_{n_{0}}^{-}]<\infty$ when $n$ goes to +$\infty$+ ∞. This contradicts the fact that $\text{cr}([q_{n}],[q_{0}])$ goes to +$\infty$+ ∞. Hence for every pair of points in $I$, the associated projective line is contained in $I$. In particular $I$ is a projective subspace. ∎
Let $\operatorname{Gr}_{d}^{(n,n)}(\mathcal{Q})$ be the open subset of $\operatorname{Gr}_{d}(\mathcal{Q})$ pf pencils $P$ such that for all $q\in P$ non-zero, $q$ has signature $(n,n)$. We now apply the previous results to prove that representations that admit an equivariant fitting immersion in this open set, and more generally a fitting map with a fitting flow, are Anosov. For examples of such immersions, see Propositon 7.1.
@@
Let $u:\widetilde{N}\to\operatorname{Gr}_{d}^{(n,n)}(\mathcal{Q})$ be a continuous $\rho$-equivariant map that admits an equivariant fitting flow. The representation $\rho$ is $\{n\}$-Anosov, and for any $q\in\mathbb{S}u^{*}\mathcal{E}$, the limit map at the limit $\zeta\in\partial\Gamma$ of the the flow line $(\Phi_{t}(q))$ can be characterized as:
$\mathbb{P}\left(\xi^{n}_{\rho}(\zeta)\right)=\bigcap_{t\geq 0}\{\pi(\Phi_{t}(q))\leq 0\}.$
Note that the flow lines of an equivariant fitting flow are quasi-geodesics by Proposition 4.5, hence the limit point $\zeta$ is well-defined.
$h^{\prime}_{q}=0$$\pi(q)=0$$h^{\prime}_{\Phi_{t}(q)}=0$$\pi(\Phi_{t}(q))=0$$e$$L$$\ell_{1}^{\prime}$$\ell_{1}$$\ell_{2}^{\prime}$$\ell_{2}$$\ell_{3}^{\prime}$$\ell_{3}$$\ell_{4}^{\prime}$$\ell_{4}$$f\text{ at }\infty\text{ on }L\text{ }\to$
@@
This defines transverse vector subspaces by Proposition 5.3 since $\text{cr}(\Phi_{t}(q),q)$ and $\text{cr}(-\Phi_{-t}(q),-q)$ go to +$\infty$+ ∞ when $t$ goes to +$\infty$+ ∞. Moreover the quadrics in the pencils in the image of $u$ are of signature $(n,n)$, one must have $\dim(E_{q})=\dim(F_{q})=n$ so this splitting is well-defined. This splitting is preserved by $\Phi$.
We now construct a metric $h$ on this flat $V$-bundle over $\mathbb{S}u^{*}\mathcal{E}$. Given $q\in\mathbb{S}u^{*}\mathcal{E}$ we define the symmetric bilinear form $h_{q}$ on $V=E_{q}\oplus F_{q}$ so that this sum is orthogonal and $h_{q}$ is equal to $\pi(q)\in\mathcal{Q}$ on $E_{q}$ and $-\pi(q)\in\mathcal{Q}$ on $F_{q}$. Note that by definition of $E_{q}$ and $F_{q}$, $h_{q}$ is positive.
We also introduce an auxiliary symmetric bilinear form $h^{\prime}$ of signature $(n,n)$ on this flat $V$-bundle over $\mathbb{S}u^{*}\mathcal{E}$ so that the sum $E_{q}\oplus F_{q}$ is orthogonal and $h^{\prime}_{q}$ is equal to $\pi(q)\in\mathcal{Q}$ on $E_{q}$ and on $F_{q}$.
Our first step is to compare the quadric $h^{\prime}_{q}$ with $\pi(q)$. Let $L$ be a projective line intersecting $\mathbb{P}(E_{q})$ at some $e$ and $\mathbb{P}(F_{q})$ at some $f$. In Figure LABEL:fig:Anosov_3, we illustrate some $\mathbb{RP}^{2}\subset\mathbb{RP}^{2n-1}$ containing the projective line $L$. Let $\ell_{1},\ell^{\prime}_{1}$ be the zeroes of $\pi(q)$ on $L$ and $\ell_{2},\ell^{\prime}_{2}$ be the zeroes of $h^{\prime}_{q}$ on $L$, so that $\ell_{1},\ell_{2}$ lie in the same connected component of $L\setminus\{e,f\}$. Since $N$ is compact, there exist a maximum $\delta<\infty$ for all $q\in\mathbb{S}u^{*}\mathcal{E}$ and all such projective line $L$ of the following quantity:
$|\log\left([e,\ell_{1},\ell_{2},f]\right)|.$
@@
Let $e,f$ be the lines generated by $v,w$ and $L$ be the projective line joining them. Let $\ell_{1},\ell^{\prime}_{1}$ be the zeroes of $\pi(q)$ on $L$, $\ell_{2},\ell^{\prime}_{2}$ the zeroes of $h^{\prime}_{q}$ on $L$, $\ell_{3},\ell^{\prime}_{3}$ be the zeroes of $\pi(\Phi_{t}(q))$ on $L$ and finally $\ell_{4},\ell^{\prime}_{4}$ the zeroes of $h_{\Phi_{t}(q)}$ on $L$. We assume that $\ell_{1},\ell_{2},\ell_{3},\ell_{4}$ all lie on the same component of $L\setminus\{e,f\}$.
The cross ratio $[e,\ell_{2},\ell_{4},f]$ is equal to $R^{\frac{1}{2}}$. Indeed $\ell_{2}$ is generated by $h^{\frac{1}{2}}_{q}(w)v+h^{\frac{1}{2}}_{q}(v)w$ and $\ell_{4}$ is generated by $h^{\frac{1}{2}}_{\Phi_{t}(q)}(w)v+h^{\frac{1}{2}}_{\Phi_{t}(q)}(v)w$, up to changing $\ell_{i}$ by $\ell_{i}^{\prime}$ for $1\leq i\leq 4$. Hence:
$[e,\ell_{2},\ell_{4},f]=\frac{\left(h^{\frac{1}{2}}_{\Phi_{t}(q)}(w)v+h^{\frac{1}{2}}_{\Phi_{t}(q)}(v)w\right)\wedge v}{\left(h^{\frac{1}{2}}_{\Phi_{t}(q)}(w)v+h^{\frac{1}{2}}_{\Phi_{t}(q)}(v)w\right)\wedge w}\times\frac{\left(h^{\frac{1}{2}}_{q}(w)v+h^{\frac{1}{2}}_{q}(v)w\right)\wedge w}{\left(h^{\frac{1}{2}}_{q}(w)v+h^{\frac{1}{2}}_{q}(v)w\right)\wedge v}=R^{\frac{1}{2}}.$
However due to our comparison of $\pi(q)$ and $h^{\prime}_{q}$ one has:
$e^{-\delta}\leq[e,\ell_{1},\ell_{2},f],[e,\ell_{3},\ell_{4},f]\leq e^{\delta}.$
Therefore $[e,\ell_{2},\ell_{4},f]/[e,\ell_{1},\ell_{3},f]\geq e^{-2\delta}$. Hence $R^{\frac{1}{2}}>e^{-2\delta}[e,\ell_{1},\ell_{3},f]$. This last cross ratio is larger than:
$[\ell_{1}^{\prime},\ell_{1},\ell_{3},\ell_{3}^{\prime}]\geq\text{cr}\left(\pi(\Phi_{t}(q)),\pi(q),\right).$
Since $\Phi$ is a fitting flow and since $N$ is compact, there exist $\alpha>0$ such that $\text{cr}\left(\pi(\Phi_{1}(q)),\pi(q),\right)\geq e^{\alpha}$ for all $q\in\mathbb{S}u^{*}\mathcal{E}$. Hence by the triangular inequality from Proposition 5.2, for all $t>0$:
$\text{cr}\left(\pi(\Phi_{t}(q)),\pi(q),\right)\geq e^{\alpha(t-1)}.$
@@ 5.2. Fibered domain of discontinuity
Proving first that $\mathbb{P}(S^{2}V^{>0})$ is fully covered first helps us proving that the Guichard -Wienhard domain is also fully covered.
@@
Let us fist prove that all of $\mathbb{P}(S^{2}V^{>0})$ is covered by the union of $\mathbb{P}(u^{\circ}(x)\cap S^{2}V^{>0})$ for $x\in\widetilde{N}$. We fix a Riemanian metric on $N$ that defines a Riemannian metric $g$ on $\widetilde{N}$, with associated distance $d_{g}$. Since $u$ is globally fitting, see Remark 4.6, for all $x\in\widetilde{N}$ there exist a neighborhood $U$ of $\mathbb{P}(u^{\circ}(x)\cap S^{2}V^{\geq 0})$ in $\mathbb{P}(S^{2}V^{\geq 0})$ that is covered by the manifolds $\mathbb{P}(u^{\circ}(y)\cap S^{2}V^{\geq 0})$ for $y$ in the ball for $d_{g}$ of radius $1$ centered at 00. This neighborhood contains an $\epsilon$-neighborhood of $\mathbb{P}(u^{\circ}(x)\cap S^{2}V^{>0})$ for the Hilbert metric on $\mathbb{P}(S^{2}V^{>0})$ for some $\epsilon>0$. Since $N$ is compact, this $\epsilon>0$ can be chosen independently of $x$.
@@
Now let us fix some $x_{0}\in\widetilde{N}$ and some $[p_{0}]\in\mathbb{P}(u^{\circ}(x_{0})\cap S^{2}V^{>0})$. Given any $[p^{\prime}]\in\mathbb{P}(S^{2}V^{>0})$ one can find a finite sequence $p_{0},p_{1},\cdots,p_{k}=p^{\prime}$ in $S^{2}V^{>0}$ so that the Hilbert distance between $[p_{i}]$ and $[p_{i+1}]$ is less than $\epsilon$ for all $0\leq i<k$. By induction, and since the Riemannian metric $d_{g}$ is complete, one can construct for all $1\leq i\leq k$ a point $x_{i}\in\widetilde{N}$ such that $[p_{i}]\in\mathbb{P}(u^{\circ}(x_{i})\cap S^{2}V^{>0})$ and $d_{H}(x_{i-1},x_{i})\leq 1$. Therefore the manifolds $\mathbb{P}(u^{\circ}(x)\cap S^{2}V^{>0})$ cover all of $\mathbb{P}(S^{2}V^{>0})$.
@@
Now let us consider the fibered domain in projective space. Consider a rank one line $[p]\in S^{2}\mathbb{P}(\xi^{n}_{\rho}(x))$ for some $\zeta\in\partial\Gamma$. Suppose that $p$ belongs to $u^{\circ}(\zeta)$ for some $x\in\widetilde{N}$. There exist a flow line $(\Phi_{t}([q]))_{t\geq 0}$ starting at $p([q])=x$ and converging to $\zeta\in\partial\Gamma\simeq\partial\widetilde{N}$ by Proposition 4.5. Theorem 5.4 implies that $\pi(q)\in\mathcal{Q}$ must be negative on $\xi^{2}_{\rho}(\zeta)$ and hence $p$ cannot belong to $u^{\circ}(x)$.
@@
Conversely fix a point $x\in\widetilde{N}$ and take any rank one point $[p]\in S^{2}\Omega$ in the Guichard-Wienhard domain of discontinuity. There exist a sequence $(x_{n})$ such that $[p]$ belongs to the limit of $\mathbb{P}(u^{\circ}(x_{n})\cap S^{2}V^{\geq 0})$, since these manifolds cover $\mathbb{P}(S^{2}V^{>0})$. We consider some $[q_{n}]\in\mathbb{S}\mathcal{E}_{x}$ such that $\Phi_{t}([q_{n}])\in\mathbb{S}\mathcal{E}_{x_{n}}$ for some $t_{n}>0$, which exist by Proposition 4.5. If $t_{n}$ diverges when $n$ varies, then the set $\mathbb{P}(u^{\circ}(x_{n})\cap S^{2}\mathbb{P}(V))$ becomes arbitrarily close to $S^{2}\mathbb{P}(\xi^{2}_{\rho}(\zeta_{n}))$ where $\zeta_{n}$ is the limit when $t$ goes to +$\infty$+ ∞ of $\Phi_{t}([q_{n}])$. This would contradict the fact that $[p]\in S^{2}\Omega$, as in this case $[p]\in\mathbb{P}(\xi_{\rho}^{n}(\zeta))$ where $\zeta$ is a limit point of $(\zeta_{n})$. Hence the sequence $(x_{n})$ is bounded and therefore converges up to subsequence to some $x_{\infty}\in\widetilde{N}$, and $[p]\in\mathbb{P}(u^{\circ}(x_{\infty}))$.
@@
Let us consider representations $\rho:\Gamma\to\operatorname{Sp}(2n,\mathbb{R})$. We prove our main result, which is the characterization of maximal representations in terms of maps of pencils. The first part introduces $\omega$-regular pencils, as well as a connected component of the space of $\omega$-regular pencils. We then state the characterization, and then present the construction of a map of pencils with a fitting flow for any maximal representation. This construction relies to a map from the space of pairs of Lagragians to the space of quadrics. We briefly discuss how one can also use this map to decompose the projective structure into polygons of quadric hypersurfaces in Section LABEL:subsec:polygons_of_hyperboloids.
@@
If a locally fitting map $u:\widetilde{S_{g}}\to\operatorname{Gr}_{2}^{{(n,n)}}(\mathcal{Q})$ admits a fitting flow which is equivariant with respect to a representation $\rho:\Gamma\to\operatorname{Sp}(2n,\mathbb{R})$, the image of $u$ must lie in $\operatorname{Gr}_{2}^{\omega}(\mathcal{Q})$ as $\rho$ is $\{n\}$-Anosov and its limit map takes values in the space of Lagrangians.
There are non-maximal representations admitting equivariant fitting immersions, for instance almost-Fuchsian representations in $\operatorname{SL}(2,\mathbb{C})\subset\operatorname{Sp}(4,\mathbb{R})$. In order to obtain the maximality property, we need to restrict ourselves to the correct union of connected component of $\operatorname{Gr}_{2}^{\omega}(\mathcal{Q})$.
@@
Let $q\in\mathcal{Q}$ be an $\omega$-regular element. The set of Lagrangians $\ell$ such that $q$ is positive on $\ell$ is an open ball.
@@
$q(v,v)+q(u(v),u(v))+2q(v,u(v))=q(v+u(v),v+u(v))>0.$
Since $q(v,v)>0$ and $q(u(v),u(v))<0$, for all $10<\lambda<$ 1:
$q(v+\lambda u(v),v+\lambda u(v))=q(v,v)+\lambda q(u(v),u(v))+2\lambda q(v,u(v))>0.$
@@
Let $P\in\operatorname{Gr}_{2}^{\omega}(\mathcal{Q})$. There exist a continuous map $\xi_{P}:\mathbb{S}P\to\mathcal{L}_{n}$ such that for all $[q]\in\mathbb{S}P$, $q>0$ on $\xi_{P}([q])$. Moreover any two such maps are homotopic, so the free homotopy type $[\xi_{P}]$ is well defined.
@@
We say that a pencil is maximal for some orientation if $[\xi_{P}]=n[\tau]$, where $[\tau]$ is the generator of $\pi_{1}(\mathcal{L}_{n})$ introduced in Section 2.2. We denote by $\operatorname{Gr}_{2}^{\max}(\mathcal{Q})$ the space of pencils that are maximal for some orientation. This is a union of connected components of $\operatorname{Gr}^{\omega}_{2}(\mathcal{Q})$ as the homotopy type $[\xi_{P}]$ is locally invariant for $P\in\operatorname{Gr}_{2}^{\omega}(\mathcal{Q})$.
@@
As a recal we have the following inclusions:
@@
Given a maximal representation $\rho:\Gamma_{g}\to\operatorname{Sp}(2n,\mathbb{R})$, is there always an equivariant fitting immersion $u:\widetilde{S_{g}}\to\operatorname{Gr}_{2}^{\max}(\mathcal{Q})$

### v2

@@ 5.1. The Anosov property
Since $z\notin I$ there exist $n_{0}\in\mathbb{N}$ such that $q_{n_{0}}$ is positive on $z$. For $n\geq n_{0}$, let $z_{1},z_{2}$ be the two intersections of $L$ with the zeroes of $q_{n}$ so that the points $(x,z_{n}^{-},z,z_{n}^{+},y)$ are cyclically ordered. The sequences $(z_{n}^{-})$ and $(z_{n}^{+})$ are monotonic in $S$ and must converge to $x$ and $y$ since $S\cap I=\emptyset$. The value of $\text{cr}([q_{n_{0}}],[q_{n}])$ is bounded from above by the cross ratio $z_{n_{0}}^{+},z_{n}^{+},z_{n}^{-},z_{n_{0}}^{-}]$, which is turns converges to the cross ratio $z_{n_{0}}^{+},x,y,z_{n_{0}}^{-}]<\infty$ when $n$ goes to $\infty$. This contradicts the fact that $\text{cr}([q_{n}],[q_{0}])$ goes to $\infty$. Hence for every pair of points in $I$, the associated projective line is contained in $I$. In particular $I$ is a projective subspace. ∎
Let $\operatorname{Gr}_{d}^{(n,n)}(\mathcal{Q})\subset\operatorname{Gr}_{d}^{\text{mix}}(\mathcal{Q})$ be the open subset of $\operatorname{Gr}_{d}(\mathcal{Q})$ of pencils $P$ such that for all $q\in P$ non-zero, $q$ has signature $(n,n)$. We now apply the previous results to prove that representations that admit an equivariant fitting immersion in this open set, and more generally a fitting map with a fitting flow, are Anosov. For examples of such immersions, see Propositon 7.1. Here again $M=\widetilde{N}$ for a compact $N$, $\Gamma=\pi_{1}(N)$ and $\rho:\Gamma\to\operatorname{SL}(V)$ is a representation.
@@
Let $u:\widetilde{N}\to\operatorname{Gr}_{d}^{(n,n)}(\mathcal{Q})$ be a continuous $\rho$-equivariant map that admits an equivariant fitting flow. The representation $\rho$ is $\{n\}$-Anosov, and for any $q\in\mathbb{S}u^{*}\mathcal{E}$, the limit map $\xi^{n}_{\rho}(\zeta)\in\operatorname{Gr}_{n}(\mathbb{R}^{2n})$ at the limit $\zeta\in\partial\Gamma$ of the the flow line $(\Phi_{t}(q))$ is the following projective subspace:
$\bigcap_{t\geq 0}\{\pi(\Phi_{t}(q))\leq 0\}.$
Note that the flow lines of an equivariant fitting flow are quasi-geodesics by Proposition 4.5, and $\mathbb{S}u^{*}\mathcal{E}$ is quasi-isometric to $\Gamma$, hence the limit point $\zeta$ is well-defined.
[…]
@@
This defines transverse vector subspaces by Proposition 5.3 since $\text{cr}(\Phi_{t}(q),q)$ and $\text{cr}(-\Phi_{-t}(q),-q)$ go to $\infty$ when $t$ goes to $\infty$. Moreover the quadrics in the pencils in the image of $u$ are of signature $(n,n)$, one must have $\dim(E_{q})=\dim(F_{q})=n$ so this splitting is well-defined. This splitting is preserved by $\Phi$.
We now construct a metric $h$ on this flat $V$-bundle over $\mathbb{S}u^{*}\mathcal{E}$. Given $q\in\mathbb{S}u^{*}\mathcal{E}$ we define the symmetric bilinear form $h_{q}$ on $V=E_{q}\oplus F_{q}$ so that this sum is orthogonal and $h_{q}$ is equal to $\pi(q)\in\mathcal{Q}$ on $E_{q}$ and $\pi(q)\in\mathcal{Q}$ on $F_{q}$. Note that by definition of $E_{q}$ and $F_{q}$, $h_{q}$ is positive.
We also introduce an auxiliary symmetric bilinear form $\bar{h}$ of signature $(n,n)$ on this flat $V$-bundle over $\mathbb{S}u^{*}\mathcal{E}$ so that the sum $E_{q}\oplus F_{q}$ is orthogonal and $\bar{h}_{q}$ is equal to $\pi(q)\in\mathcal{Q}$ on $E_{q}$ and on $F_{q}$.
Our first step is to compare the quadric $\bar{h}_{q}$ with $\pi(q)$. Let $L$ be a projective line intersecting $\mathbb{P}(E_{q})$ at some $e$ and $\mathbb{P}(F_{q})$ at some $f$. In Figure 3, we illustrate some $\mathbb{RP}^{2}\subset\mathbb{RP}^{2n-1}$ containing the projective line $L$. Let $\ell_{1},\ell^{\prime}_{1}$ be the zeroes of $\pi(q)$ on $L$ and $\ell_{2},\ell^{\prime}_{2}$ be the zeroes of $\bar{h}_{q}$ on $L$, so that $\ell_{1},\ell_{2}$ lie in the same connected component of $L\setminus\{e,f\}$. Since $N$ is compact, there exist a maximum $\delta<\infty$ for all $q\in\mathbb{S}u^{*}\mathcal{E}$ and all such projective line $L$ of the quantity $|\log\left([\ell_{1},\ell_{2},f,e]\right)|$.
@@
Let $e,f$ be the lines generated by $v,w$ and $L$ be the projective line joining them. Let $\ell_{1},\ell^{\prime}_{1}$ be the zeroes of $\pi(q)$ on $L$, $\ell_{2},\ell^{\prime}_{2}$ the zeroes of $\bar{h}_{q}$ on $L$, $\ell_{3},\ell^{\prime}_{3}$ be the zeroes of $\pi(\Phi_{t}(q))$ on $L$ and finally $\ell_{4},\ell^{\prime}_{4}$ the zeroes of $h_{\Phi_{t}(q)}$ on $L$. We assume that $\ell_{1},\ell_{2},\ell_{3},\ell_{4}$ all lie on the same component of $L\setminus\{e,f\}$.
The cross ratio $\ell_{2},\ell_{4},f,e]$ is equal to $R^{\frac{1}{2}}$. Indeed $\ell_{2}$ is generated by $h^{\frac{1}{2}}_{q}(w)v+h^{\frac{1}{2}}_{q}(v)w$ and $\ell_{4}$ is generated by $h^{\frac{1}{2}}_{\Phi_{t}(q)}(w)v+h^{\frac{1}{2}}_{\Phi_{t}(q)}(v)w$, up to changing $\ell_{i}$ by $\ell_{i}^{\prime}$ for $1\leq i\leq 4$. Hence:
$\ell_{2},\ell_{4},f,e]=\frac{\left(h^{\frac{1}{2}}_{q}(w)v+h^{\frac{1}{2}}_{q}(v)w\right)\wedge w}{\left(h^{\frac{1}{2}}_{q}(w)v+h^{\frac{1}{2}}_{q}(v)w\right)\wedge v}\times\frac{\left(h^{\frac{1}{2}}_{\Phi_{t}(q)}(w)v+h^{\frac{1}{2}}_{\Phi_{t}(q)}(v)w\right)\wedge v}{\left(h^{\frac{1}{2}}_{\Phi_{t}(q)}(w)v+h^{\frac{1}{2}}_{\Phi_{t}(q)}(v)w\right)\wedge w}=R^{\frac{1}{2}}.$
However due to our comparison of $\pi(q)$ and $\bar{h}_{q}$ one has:
$e^{-\delta}\leq[\ell_{2},\ell_{4},f,e],[\ell_{1},\ell_{3},f,e]\leq e^{\delta}.$
Therefore $\ell_{2},\ell_{4},f,e]/[\ell_{1},\ell_{3},f,e]\geq e^{-2\delta}$. Hence $R^{\frac{1}{2}}>e^{-2\delta}[\ell_{1},\ell_{3},f,e]$. This last cross ratio is larger than:
$\ell_{1},\ell_{3},\ell_{3}^{\prime},\ell_{1}^{\prime}]\geq\text{cr}\left(\pi(\Phi_{t}(q)),\pi(q),\right).$
Since $\Phi$ is a fitting flow and since $N$ is compact, there exist $\alpha>0$ such that $\text{cr}\left(\pi(\Phi_{1}(q)),\pi(q)\right)\geq e^{\alpha}$ for all $q\in\mathbb{S}u^{*}\mathcal{E}$. Hence by the triangular inequality from Proposition 5.2, for all $t>0$:
$\text{cr}\left(\pi(\Phi_{t}(q)),\pi(q)\right)\geq e^{\alpha(t-1)}.$
@@ 5.2. Fibered domain of discontinuity
Recall that $u^{o}(x)\subset S^{2}V$ is the anihilator subspace of $u(x)\subset\mathcal{Q}=\left(S^{2}V\right)^{*}$.
@@
Proving first that $\mathbb{P}(S^{2}V^{>0})$ is fully covered helps us proving that the Guichard -Wienhard domain is also fully covered.
@@
Let us fist prove that all of $\mathbb{P}(S^{2}V^{>0})$ is covered by the union of $\mathbb{P}(u^{\circ}(x)\cap S^{2}V^{>0})$ for $x\in\widetilde{N}$. We fix a Riemanian metric on $N$ that defines a Riemannian metric $g$ on $\widetilde{N}$, with associated distance $d_{g}$. Since $u$ is globally fitting, see Remark 4.6, for all $x\in\widetilde{N}$ there exist a neighborhood $U$ of $\mathbb{P}(u^{\circ}(x)\cap S^{2}V^{\geq 0})$ in $\mathbb{P}(S^{2}V^{\geq 0})$ that is covered by the manifolds $\mathbb{P}(u^{\circ}(y)\cap S^{2}V^{\geq 0})$ for $y$ in the ball for $d_{g}$ of radius $1$ centered at 0, by Proposition 3.12, part (i).
This neighborhood contains an $\epsilon$-neighborhood of $\mathbb{P}(u^{\circ}(x)\cap S^{2}V^{>0})$ for the Hilbert metric on $\mathbb{P}(S^{2}V^{>0})$ for some $\epsilon>0$. Since $N$ is compact, this $\epsilon>0$ can be chosen independently of $x$.
@@
Now let us fix some $x_{0}\in\widetilde{N}$ and some $s_{0}]\in\mathbb{P}(u^{\circ}(x_{0})\cap S^{2}V^{>0})$. Given any $s^{\prime}]\in\mathbb{P}(S^{2}V^{>0})$ one can find a finite sequence $s_{0},s_{1},\cdots,s_{k}=s^{\prime}$ in $S^{2}V^{>0}$ so that the Hilbert distance between $s_{i}]$ and $s_{i+1}]$ is less than $\epsilon$ for all $0\leq i<k$. By induction, and since the Riemannian metric $d_{g}$ is complete, one can construct for all $1\leq i\leq k$ a point $x_{i}\in\widetilde{N}$ such that $s_{i}]\in\mathbb{P}(u^{\circ}(x_{i})\cap S^{2}V^{>0})$ and $d_{H}(x_{i-1},x_{i})\leq 1$. Therefore the manifolds $\mathbb{P}(u^{\circ}(x)\cap S^{2}V^{>0})$ cover all of $\mathbb{P}(S^{2}V^{>0})$.
@@
Now let us consider the fibered domain in projective space. We first prove that the fibers are contained in the domain $S^{2}\Omega$. Consider a rank one line $s]\in S^{2}\mathbb{P}(\xi^{n}_{\rho}(x))$ for some $\zeta\in\partial\Gamma$. Suppose that $p$ belongs to $u^{\circ}(x)$ for some $x\in\widetilde{N}$. There exist a flow line $(\Phi_{t}([q]))_{t\geq 0}$ starting at $\operatorname{pr}([q])=x$ and converging to $\zeta\in\partial\Gamma\simeq\partial\widetilde{N}$ by Proposition 4.5. Theorem 5.4 implies that $\pi(q)\in\mathcal{Q}$ must be negative on $\xi^{2}_{\rho}(\zeta)$ and hence $p$ cannot belong to $u^{\circ}(x)$.
@@
Conversely let us prove that every point in $S^{2}\Omega$ is covered by some fiber. Fix a point $x\in\widetilde{N}$ and take any rank one point $s]\in S^{2}\Omega$ in the Guichard-Wienhard domain of discontinuity. There exist a sequence $(x_{n})$ such that $[s]$ belongs to the limit of $\mathbb{P}(u^{\circ}(x_{n})\cap S^{2}V^{\geq 0})$, since these manifolds cover $\mathbb{P}(S^{2}V^{>0})$. We consider some $q_{n}]\in\mathbb{S}\mathcal{E}_{x}$ such that $\Phi_{t}([q_{n}])\in\mathbb{S}\mathcal{E}_{x_{n}}$ for some $t_{n}>0$, which exist by Proposition 4.5. If $t_{n}$ diverges when $n$ varies, then the set $\mathbb{P}(u^{\circ}(x_{n})\cap S^{2}\mathbb{P}(V))$ becomes arbitrarily close to $S^{2}\mathbb{P}(\xi^{n}_{\rho}(\zeta_{n}))$ where $\zeta_{n}$ is the limit when $t$ goes to $\infty$ of $\Phi_{t}([q_{n}])$. This would contradict the fact that $s]\in S^{2}\Omega$, as in this case $s]\in\mathbb{P}(\xi_{\rho}^{n}(\zeta))$ where $\zeta$ is a limit point of $(\zeta_{n})$. Hence the sequence $(x_{n})$ is bounded and therefore converges up to subsequence to some $x_{\infty}\in\widetilde{N}$, and $s]\in\mathbb{P}(u^{\circ}(x_{\infty}))$.
@@
Let us consider representations $\rho:\Gamma\to\operatorname{Sp}(2n,\mathbb{R})$. We prove our main result, which is the characterization of maximal representations in terms of maps of pencils. The first part introduces $\omega$-regular pencils, as well as a connected component of the space of $\omega$-regular pencils. We then state the characterization, and then present the construction of a map of pencils with a fitting flow for any maximal representation. This construction relies to a map from the space of pairs of Lagragians to the space of quadrics.
@@
If a locally fitting map $u:\widetilde{S_{g}}\to\operatorname{Gr}_{2}^{{(n,n)}}(\mathcal{Q})$ admits a fitting flow which is equivariant with respect to a representation $\rho:\Gamma\to\operatorname{Sp}(2n,\mathbb{R})$, the image of $u$ must lie in $\operatorname{Gr}_{2}^{\omega}(\mathcal{Q})$ as $\rho$ is $\{n\}$-Anosov by Theorem 5.4, and the limit map of $\{n\}$-Anosov representations in $\operatorname{Sp}(2n,\mathbb{R})$ take values in the space of Lagrangians. This last fact is a consequence of the fact that an attractive fixed $n$-dimensional subspace of an element in $\operatorname{Sp}(2n,\mathbb{R})$ is necessarily Lagrangian, see for instance [GW12].
There are non-maximal representations admitting equivariant fitting immersions, for instance almost-Fuchsian representations in $\operatorname{SL}(2,\mathbb{C})\subset\operatorname{Sp}(4,\mathbb{R})$, see Appendix A. In order to obtain the maximality property, we need to restrict ourselves to the correct union of connected component of $\operatorname{Gr}_{2}^{\omega}(\mathcal{Q})$.
@@
Let $q\in\mathcal{Q}$ be an $\omega$-regular element. The set of Lagrangians $\ell$ such that $q$ is positive on $\ell$ is homeomorphic to an open ball.
@@
q(v,v)+q(u(v),u(v))+2q(v,u(v))=q(v+u(v),v+u(v))>0.q(v,v)+q(u(v),u(v))+2q(v,u(v))=q(v+u(v),v+u(v))>0.
Since $q(v,v)>0$ and $q(u(v),u(v))<0$, for all $0<\lambda<1$:
$q(v+\lambda u(v),v+\lambda u(v))=q(v,v)+\lambda^{2}q(u(v),u(v))+2\lambda q(v,u(v))>0.$
@@
Let $P\in\operatorname{Gr}_{2}^{\omega}(\mathcal{Q})$. There exist a continuous map $\xi_{P}:\mathbb{S}P\to\mathcal{L}_{n}$ such that for all $q]\in\mathbb{S}P$, $q>0$ on $\xi_{P}([q])$. Moreover any two such maps are homotopic, so the free homotopy type $\xi_{P}]$ is well defined.
@@
We say that a pencil is maximal for some orientation if $\xi_{P}]=n[\tau]$, where $\tau]$ is the generator of $\pi_{1}(\mathcal{L}_{n})$ introduced in Section 2.2. We denote by $\operatorname{Gr}_{2}^{\max}(\mathcal{Q})$ the space of pencils that are maximal for some orientation. This is a union of connected components of $\operatorname{Gr}^{\omega}_{2}(\mathcal{Q})$ as the homotopy type $\xi_{P}]$ is locally invariant for $P\in\operatorname{Gr}_{2}^{\omega}(\mathcal{Q})$.
@@
As a recall we have the following inclusions:
@@
Given a maximal representation $\rho:\Gamma_{g}\to\operatorname{Sp}(2n,\mathbb{R})$, is there always an equivariant fitting immersion $u:\widetilde{S_{g}}\to\operatorname{Gr}_{2}^{\max}(\mathcal{Q})$ ?
@@
Recall that the limit map of $\{n\}$-Anosov representations in $\operatorname{Sp}(2n,\mathbb{R})$ take values in the space of Lagrangians, see Remark 6.1.